Weekly rate ($1,827 x 12)/52-$421.615 rounded to $421.62 Hourly rate $421.62/40 $10.540 rounded to $10.54 OT. rate $10.54 x 1.5 $15.81 se the minimum hourly wage of $7,25 in solving these problems eand all that follow Example 2-7 Joanne French works a 40-hour week at S 12.40 an hour with overtime hours paid at 1 ½ times the regular rate Regular weekly earnings are The overtime rate is Ir French works 4 hours overtime, additional earnings for the 4 hours are Frenchs weekly gross earnings are If paid only for time actually worked and she works only 36 hours during a week, French earns $496.00 (40 x $12.40) $18.60 ($12.40 x 1.5) $74.40 (4 × $18.60) $570.40 ($496.00+$74.40) $446.40 (36 x $12.40) The hours worked and the hourly wage rates for five employees of Koogel Company for the week ended June 10 follow. If required, round your answers to the nearest cent. a. For each employee, compute the gross earnings b. Compute the total gross earnings for all employees Employee Duffy, M. Hazelton, G. Inman, T. Palmer, C Diaz, O. Hours Worked 35 28 39h 24% 31 Regular Hourly Wage Rate $11.90 14.15 10.90 16.40 12.90 (b) Total gross earnings Gross Earnings Check My Work All work saved.
Answer #1
Employee Hours Worked Regular Hourly Wage Rate Gross Earnings
Du±y, M. 35 11.90 416.50 =35*11.9
Hazelton, G. 28 14.15 396.20 =28.14.15
Inman, T. 39 1/2 10.90 430.55 =39.5*10.9
Palmer, C. 24 1/2 16.40 401.80 =24.5*16.4
Diaz, O. 31 12.90 399.90 =31*12.9
Total gross earnings 2044.95